Operations Report Dec 31st

Crew 202 Operations Report 31-DEC-2018

SOL: 2

Name of person filing report: Kasey Hilton

Non-nominal systems: Suits #3 & #7, electric tea kettle, and hot water heater

Notes on non-nominal systems: Two-piece suit #3 and one-piece suit #7 won’t turn on even when plugged in; Electric tea kettle in kitchen does not turn on even when we plugged it into a different outlet; Hot water heater ignitor doesn’t work so there is not hot water in the Hab.

Generator (hours run): 7hr 47min; Turned off around 9:38 AM; Turned on at 5:25 PM

Solar – SOC 86%

Diesel Reading – 75%

Propane Reading – 57%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – Full; 530 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – Full; 280 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 6:24 PM to refill loft tank

Water Meter: 01397146 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 63.4

Beginning charge: 84%

Ending charge: 74%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 45.2

Beginning charge: 90%

Ending charge: 83%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 61.5

Beginning charge: 100%

Ending charge: 74%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Honda, YODA (300), 350.1 (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Training for geology EVA

Oil Added? No

ATV Fuel Used: Tanks are still full; Only used for about 10 minutes each

# Hours the ATVs were Used today: 1hr

Notes on ATVs: All VIN numbers on ATVs were located and registration papers were found for one. YODA, Honda, and two of the 350s were started and driven today, but they were very hard to start in the cold. The Honda, YODA, and 350.1 remain at the Hab for a geology EVA later in the rotation. There are two other 350s that are being worked on since one battery doesn’t work and the other is making a weird noise. One of these 350s will be brought to the Hab later.

HabCar used and why, where? Off site

CrewCar used and why, where? Off site

General notes and comments: Spent the first half of the day training for EVAs and ATVs. The second half of the day was spend doing two EVAs, both to Marble Ritual and back.

Summary of internet: Nothing to report

Summary of suits and radios: Two-piece suit #3 and one piece suit #7 air flow not working as reported from yesterday and mentioned above in “Non-nominal systems”

Summary of Hab operations: The hot water heater igniter doesn’t work and there is not hot water in the Hab as mentioned above in “Non-nominal systems”

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: I have not heard anything about the two-piece suit #3 since it was reported about 2 days ago.